solve for proportions
52 over 4 equals m over 5

Answer by jim_thompson5910(35256)

52/4 = m/5

52*5 = 4m

260 = 4m

4m = 260

m = 260/4

m = 65

Answer: m = 65

--------------------------------------------------------------------------------------------------------------

If you need more help, you can email me at jim_thompson5910@hotmail.com

To find other ways to contact me, you can also visit my website: http://www.freewebs.com/jimthompson5910/home.html

Thanks,

Jim

--------------------------------------------------------------------------------------------------------------